OpenStudy (anonymous):

Which set of values contain three perfect squares? {1, 2, 3} {5, 25, 125} {10, 50, 100} {16, 81, 121}

OpenStudy (anonymous):

D. 4*4=16 9*9=81 11*11=121

OpenStudy (anonymous):

yes answer is D